# ExCyTIn-Bench: Evaluating LLM agents on Cyber Threat Investigation

### v1

@@ Abstract
We present ExCyTIn-Bench, the first benchmark to Evaluate an LLM agent x on the task of Cyber Threat Investigation through security questions derived from investigation graphs. Real-world security analysts must sift through a large number of heterogeneous alert signals and security logs, follow multi-hop chains of evidence, and compile an incident report. With the developments of LLMs, building LLM-based agents for automatic thread investigation is a promising direction. To assist the development and evaluation of LLM agents, we construct a dataset from a controlled Azure tenant that covers 8 simulated real-world multi-step attacks, 57 log tables from Microsoft Sentinel and related services, and 589 automatically generated questions. We leverage security logs extracted with expert-crafted detection logic to build threat investigation graphs, and then generate questions with LLMs using paired nodes on the graph, taking the start node as background context and the end node as answer. Anchoring each question to these explicit nodes and edges not only provides automatic, explainable ground truth answers but also makes the pipeline reusable and readily extensible to new logs. This also enables the automatic generation of procedural tasks with verifiable rewards, which can be naturally extended to training agents via reinforcement learning. Our comprehensive experiments with different models confirm the difficulty of the task: with the base setting, the average reward across all evaluated models is 0.249, and the best achieved is 0.368, leaving substantial headroom for future research. Code and data are coming soon!
@@ 1 Introduction
The growing reliance on digital services for critical functions worldwide underscores the need to secure our digital future. Meanwhile, cyberattacks are growing in quantity, variety, and sophistication. For example, cloud environment intrusions increased by 75% from 2022 to 2023 [8]. Although traditional defenses like behavioral analysis, malware signature matching, and anomaly detection can mitigate threats [15, 12], attackers continue to develop tactics to evade them [31]. Thus, human-led threat investigations have become critical, requiring analysts to manually go through system and network logs, apply reasoning, and leverage domain expertise to detect and respond to threats [2].
Meanwhile, advancement of Large Language Models (LLMs) has enabled astonishing achievements in complex tasks [54, 14, 45, 47, 46], that LLM agents can understand observations and select actions in complex environments such as code interpretation and database interaction to perform sequential actions [53, 63, 50, 49, 25]. Also, LLMs trained with enormous corpora of text provide them a wealth of knowledge across a range of domains [61], including cybersecurity knowledge. Thus, Cyber-Security threat investigation is a promising area for the application of LLM-based autonomous agents, as previous works have shown that LLMs are capable of multi-step observation, reasoning, and actions, which are key components for successful investigation and detection of potential threat actors and indicator of compromises (IoCs) [13, 23, 62].
@@
To assist development, a rigorous, standardized benchmark is needed to evaluate the cybersecurity investigation capability of LLM agents. The benchmark should resemble the real-world threat investigation scenario with a critical mass of security event logs, heterogeneity across real-world multi-stage security incident types. However, existing literature focuses on evaluating LLMs from a knowledge memorization perspective [56, 22], instead of targeting the security investigation and reasoning ability of LLM agents. For example, CTIBench [1] constructs a multi-choice Q&A to evaluate LLMs on Cyber Threat Intelligence(CTI) knowledge, SECURE [5] evaluates LLMs on security extraction and understanding.
To fill this gap, we build ExCyTIn-Bench, to evaluate LLM agents on cybersecurity threat investigation. Construction of the benchmark consisted of three steps: 1. Data Collection. From a fictional Microsoft Azure tenant, we collect 59 distinct tables of logs consisting of 8 different cyberattacks. 2. Question Generation. We propose a principled method to construct bipartite incident graphs from the alerts and entities involved in the attacks, then use it to generate QA pairs, resulting in a test dataset of 589 questions. 3. Environment Construction. We construct a MySQL Docker image to be a reinforcement learning (RL) environment in which agents can submit queries and receive feedback similar to InterCode [53]. The database queries are treated as actions, and the execution result as observations. We use an LLM as an evaluator by default, but deterministic checking of the answer is also available. Since our questions are generated from paths in the incident graphs, we can assign partial rewards if the agents find any intermediate information along the path.
We test with a wide range of current language models, including proprietary and open-source models, and models with different sizes and types (chat, reasoning, etc) in Section 4. We include a detailed analysis from the perspective of performance, behavior, and efficiency. We found that our benchmark is challenging even among the latest, highest-performing models, with o4-mini achieving the highest reward of 0.368. We also test different methods (e.g., ReAct [55], Expel [64], Best-of-N, Self-Reflection [38]) to help understand how different prompting and test-time scaling strategies affect performance on our benchmark. To summarize, our contributions are the following:
@@ 2.1 Background
Cybersecurity threat investigation systematically probes digital environments to detect, analyze, and mitigate malicious activity [18, 7]. Threat analysts serve as cyber-detectives: they parse vast logs, link evidence from diverse sources, and judge potential threats. Since modern infrastructures produce vast volumes of data daily, a central challenge is efficiently identifying and correlating threat signals, with other key skills required (e.g., evidence gathering and synthesis, more in Appendix B). Our benchmark is designed to assess LLM agents on all of these capabilities.
@@ 2.2 Related Work
LLMs have become a promising foundation to build agents for various complicated tasks [50, 64, 14, 6, 59]. Our goal is to create a cybersecurity threat investigation benchmark for LLM agents, which is closely related to LLM-based agents and cybersecurity: this requires LLM agents to have cybersecurity domain expertise and knowledge to be able to explore system logs, analyze suspicious behavior, and answer security-related questions. On the other hand, we build a SQL environment for LLM agents to interact with, to test models’ ability on effective and efficient SQL generation.
@@
LLMs in Cybersecurity. Most recent LLM-based cybersecurity efforts focus on knowledge memorization and information extraction [30, 24, 29, 40, 56]. CTIBench [1] evaluates LLMs’ understanding of the threat landscape via the MITRE ATT&CK framework [32], while Crimson [21] fine-tunes an LLM to map CVEs to MITRE ATT&CK techniques and generate actionable insights. SECURE [5] benchmarks models on security extraction, comprehension, and reasoning. [39] aggregates public CTI reports with structured intelligence, and [11] extracts threat behaviors from unstructured OSCTI text. Perrina et al.’s tool [35] produces CTI reports from entity graphs, and [37] uses reinforcement learning to simulate LLM-driven attacks on network topologies. Finally, CyBench [60] focuses on capture-the-flag (CTF) tasks. Although some prior work also employs graphs [11, 15], our graph-based approach differs in both concept and construction.
@@
Benchmark LLMs in Interactive Environments. [16] benchmarks LLM-based agents on data analysis tasks through an execution environment. [33] introduces a database question answering system that LLMs need to interact with a SQL interpreter, reason, and organize the results. [53] creates interactive code environments (Bash, SQL, and Python) based on static datasets [57, 28, 4] for LLM to act on. [17] build a dataset of machine learning tasks that LLMs need to perform actions like reading/writing files, executing code. SWE-Bench [20] builds a dataset on real-world software engineering problems. [52] introduces a benchmark that supports cross-environment GUI tasks over websites, desktop computers, or mobile phones.
@@
LLMs in Text-to-SQL Text-to-SQL benchmarks [57, 26] are proposed to test models on generating SQL queries given a question. Many works have been proposed to solve this task [3, 41]. [44] proposes a multi-agent framework, including a decomposer agent for Text-to-SQL generation and two auxiliary agents for tool execution and error refinement. [10] provides a systematic review of prompt engineering for Text-to-SQL generation. C3-SQL [9], StructGPT [19], Din-SQL [36] propose frameworks targeting SQL generation that consist of several stages with different strategies such as self-consistency [48] or query decomposition. StateFlow [51] introduces a framework with state and transitions to control the data exploration and selection in SQL tasks.
@@ 3.2 Question Generation
We want to create questions that can measure LLM investigation skill: to answer a question, the LLM agent needs to probe into the log data, analyze, and link related events to find the source of the malicious activity. While manual creation of these questions is expensive and does not scale, LLMs have proven useful in QA generation [5, 1]. A straightforward way is to ask an LLM to read all the details of an incident and draft Q&As. However, we found that this method produced generic questions that ignore the key concepts used to link the alerts, events, and entities together into a cohesive investigation. Moreover, these questions can lack a deterministic answer or ask about security knowledge not present in the database.
@@
4:     $V_{s}\leftarrow\textsc{Sample}(\textsc{GetFarthestEntities}(u_{s},u_{t}),k)$triangleright Get start entities
5:     $v_{e}\leftarrow\textsc{Sample}(\textsc{GetFarthestEntities}(u_{t},u_{s}),1)$triangleright Get answer
6:     $(q,a)\leftarrow M(P_{g},u_{s},V_{s},u_{t},v_{e})$triangleright Generate question and answer
7:     $s\leftarrow M(P_{s},q,\textsc{ShortestPath}(u_{s},v_{e}))$triangleright Generate solution path
@@ 3.3 Environment Setup
[…]
4:     return $1$triangleright Perfect match on final answer
@@
[…]
@@
To conduct a security investigation, the agent needs to interact with a given database. We set up a MYSQL docker environment following  [53] to execute queries. In the environment, the agent can choose to output a query to be executed or submit the answer. We set a maximum number of steps to answer one question, so the solving process ends when the agent submits the answer or the maximum number of steps is reached. We set the maximum number of entries and the maximum character length that can be returned. Only when the returned content exceeds both the max character length and the max number of entries, do we cut the entries to satisfy the max number of entries.
By default, we employ LLM to evaluate the submitted answers (we use GPT-4o). Our evaluation consists of two steps (See Algorithm 2): 1. match the submitted answer $a_{sub}$ with the ground-truth. 2. If not correct, we check if the submitted answer contains any intermediate step solutions, which shows how much progress the agent makes and useful information it acquires. We assign a decayed reward starting from the last step. This reward not only checks the final answer but also performs a fine-grained evaluation of the agent’s intermediate steps. This also evaluates the agent’s ability to identify and extract key information (IoCs) from its investigation process. Since each step of the solution contains an unambiguous string, deterministic string comparison is allowed. Thus, the environment can provide reliable rewards for intermediate steps, which is a source of process supervision that is rarely available in other reasoning tasks such as math or coding [27]. This makes our environment particularly well-suited for future work on training agents with RL.
@@ 4.2 Comparison of Different Methods
Setup. To disentangle how different methods (e.g, prompting, test-time scaling) influence agent behavior, we evaluate six different configurations (Prompts in Figure 10 to 13). (1) Base The default prompt. (2) Strategy adds additional notes on using the alert tables as a reference to investigate, which is usually how a SOC analysts find information. (3) ReAct We follow ReAct [55] to add 3 few-shot examples to base. These examples are selected from successful runs of a sampled train set. (4) Expel  [64] first distills rules from a small training split, then invokes this external memory, as well as retrieving similar examples during inference. (5) Best-of-N retries at most three times and returns the highest-reward trajectory. This is an oracle setting in which the agent needs the reward to determine if a problem needs to be rerun. (6) Reflection [38] extends Best-of-$N$ by letting the agent criticize its failed attempt, append the learned rule to the prompt, and retry. We run all methods on GPT-4o, GPT-4o-mini, and o3-mini (temp = 0, max_step = 15) and report average reward, interaction turns, and API cost in Table 3. (Also see a preliminary fine-tuning result on GPT-4o in Appendix E.2.)
@@ 5 Conclusion
Future Directions. Open-source models are catching up with propriety models in our environment, and they can be used for further distillation and training to boost performance. Our environment offers fine-grained process rewards that allow precise credit assignment across steps, which is an uncommon but valuable feature for our environment. Overall, this makes it a promising testbed for training LLM agents with reinforcement learning [65]. Another interesting direction is to utilize the inherent graph structure of the environment to define alternate state-action spaces for RL training.
@@ Appendix B Additional Background
  • Leveraging Cyber-security Domain Expertise: Applying deep knowledge of security frameworks and best practices (e.g., MITRE ATT&CK  [43], CVE (Common Vulnerabilities and Exposures) [42], OSINT (Open Source Intelligence) [34], etc.) to guide investigation processes and validate findings. Drawing on historical context about common threat actor tactics and industry-specific threats to anticipate potential entry points or attack vectors.

### v2

@@ Abstract
We present ExCyTIn-Bench, the first benchmark to Evaluate an LLM agent x on the task of Cyber Threat Investigation through security questions derived from investigation graphs. Real-world security analysts must sift through a large number of heterogeneous alert signals and security logs, follow multi-hop chains of evidence, and compile an incident report. With the developments of LLMs, building LLM-based agents for automatic thread investigation is a promising direction. To assist the development and evaluation of LLM agents, we construct a dataset from a controlled Azure tenant that covers 8 simulated real-world multi-step attacks, 57 log tables from Microsoft Sentinel and related services, and 589 automatically generated questions. We leverage security logs extracted with expert-crafted detection logic to build threat investigation graphs, and then generate questions with LLMs using paired nodes on the graph, taking the start node as background context and the end node as answer. Anchoring each question to these explicit nodes and edges not only provides automatic, explainable ground truth answers but also makes the pipeline reusable and readily extensible to new logs. This also enables the automatic generation of procedural tasks with verifiable rewards, which can be naturally extended to training agents via reinforcement learning. Our comprehensive experiments with different models confirm the difficulty of the task: with the base setting, the average reward across all evaluated models is 0.249, and the best achieved is 0.368, leaving substantial headroom for future research. The code is available at SecRL.
@@ 1 Introduction
The growing reliance on digital services for critical functions worldwide underscores the need to secure our digital future. Meanwhile, cyberattacks are growing in quantity, variety, and sophistication. For example, cloud environment intrusions increased by 75% from 2022 to 2023 (CrowdStrike, 2024). Although traditional defenses like behavioral analysis, malware signature matching, and anomaly detection can mitigate threats (Hossain et al., 2017; Hassan et al., 2020), attackers continue to develop tactics to evade them (Mahboubi et al., 2024). Thus, human-led threat investigations have become critical, requiring analysts to manually go through system and network logs, apply reasoning, and leverage domain expertise to detect and respond to threats (Alevizos and Dekker, 2024).
Meanwhile, advancement of Large Language Models (LLMs) has enabled astonishing achievements in complex tasks (Yang et al., 2024; Hong et al., 2023; Wang et al., 2023a, b, 2024b), that LLM agents can understand observations and select actions in complex environments such as code interpretation and database interaction to perform sequential actions (Yang et al., 2023; Zhang et al., 2023; Wu et al., 2024a, 2023; Li et al., 2023b). Also, LLMs trained with enormous corpora of text provide them a wealth of knowledge across a range of domains (Zhang et al., 2024b), including cybersecurity knowledge. Thus, Cyber-Security threat investigation is a promising area for the application of LLM-based autonomous agents, as previous works have shown that LLMs are capable of multi-step observation, reasoning, and actions, which are key components for successful investigation and detection of potential threat actors and indicator of compromises (IoCs) (Hassanin and Moustafa, 2024; Levi et al., 2025; Zhang et al., 2025).
@@
To assist development, a rigorous, standardized benchmark is needed to evaluate the cybersecurity investigation capability of LLM agents. The benchmark should resemble the real-world threat investigation scenario with a critical mass of security event logs, heterogeneity across real-world multi-stage security incident types. However, existing literature focuses on evaluating LLMs from a knowledge memorization perspective (Yong et al., 2025; Jing et al., 2024), instead of targeting the security investigation and reasoning ability of LLM agents. For example, CTIBench (Alam et al., 2024) constructs a multi-choice Q&A to evaluate LLMs on Cyber Threat Intelligence(CTI) knowledge, SECURE (Bhusal et al., 2024) evaluates LLMs on security extraction and understanding.
To fill this gap, we build ExCyTIn-Bench, to evaluate LLM agents on cybersecurity threat investigation. Construction of the benchmark consisted of three steps: 1. Data Collection. From a fictional Microsoft Azure tenant, we collect 59 distinct tables of logs consisting of 8 different cyberattacks. 2. Question Generation. We propose a principled method to construct bipartite incident graphs from the alerts and entities involved in the attacks, then use it to generate QA pairs, resulting in a test dataset of 589 questions. 3. Environment Construction. We construct a MySQL Docker image to be a reinforcement learning (RL) environment in which agents can submit queries and receive feedback similar to InterCode (Yang et al., 2023). The database queries are treated as actions, and the execution result as observations. We use an LLM as an evaluator by default, but deterministic checking of the answer is also available. Since our questions are generated from paths in the incident graphs, we can assign partial rewards if the agents find any intermediate information along the path.
We test with a wide range of current language models, including proprietary and open-source models, and models with different sizes and types (chat, reasoning, etc) in Section 4. We include a detailed analysis from the perspective of performance, behavior, and efficiency. We found that our benchmark is challenging even among the latest, highest-performing models, with o4-mini achieving the highest reward of 0.368. We also test different methods (e.g., ReAct (Yao et al., 2022), Expel (Zhao et al., 2024), Best-of-N, Self-Reflection (Shinn et al., 2023)) to help understand how different prompting and test-time scaling strategies affect performance on our benchmark. To summarize, our contributions are the following:
@@ 2.1 Background
Cybersecurity threat investigation systematically probes digital environments to detect, analyze, and mitigate malicious activity (IBM, 2023; CrowdStrike, 2023). Threat analysts serve as cyber-detectives: they parse vast logs, link evidence from diverse sources, and judge potential threats. Since modern infrastructures produce vast volumes of data daily, a central challenge is efficiently identifying and correlating threat signals, with other key skills required (e.g., evidence gathering and synthesis, more in Appendix B). Our benchmark is designed to assess LLM agents on all of these capabilities.
@@ 2.2 Related Work
LLMs have become a promising foundation to build agents for various complicated tasks (Wu et al., 2024a; Zhao et al., 2024; Hong et al., 2023; Chen et al., 2023; Zeng et al., 2024). Our goal is to create a cybersecurity threat investigation benchmark for LLM agents, which is closely related to LLM-based agents and cybersecurity: this requires LLM agents to have cybersecurity domain expertise and knowledge to be able to explore system logs, analyze suspicious behavior, and answer security-related questions. On the other hand, we build a SQL environment for LLM agents to interact with, to test models’ ability on effective and efficient SQL generation.
@@
LLMs in Cybersecurity. Most recent LLM-based cybersecurity efforts focus on knowledge memorization and information extraction (Liu et al., 2024; Li et al., 2023a; Liu, 2023; Sultana et al., 2023; Yong et al., 2025). CTIBench (Alam et al., 2024) evaluates LLMs’ understanding of the threat landscape via the MITRE ATT&CK framework (MITRE, 2025), while Crimson (Jin et al., 2024) fine-tunes an LLM to map CVEs to MITRE ATT&CK techniques and generate actionable insights. SECURE (Bhusal et al., 2024) benchmarks models on security extraction, comprehension, and reasoning. Siracusano et al. (2023) aggregates public CTI reports with structured intelligence, and Gao et al. (2021) extracts threat behaviors from unstructured OSCTI text. Perrina et al.’s tool (Perrina et al., 2023) produces CTI reports from entity graphs, and Rigaki et al. (2023) uses reinforcement learning to simulate LLM-driven attacks on network topologies. Finally, CyBench (Zhang et al., 2024a) focuses on capture-the-flag (CTF) tasks. Although some prior work also employs graphs (Gao et al., 2021; Hossain et al., 2017), our graph-based approach differs in both concept and construction.
@@
Benchmark LLMs in Interactive Environments. Hu et al. (2024) benchmarks LLM-based agents on data analysis tasks through an execution environment. Nan et al. (2023) introduces a database question answering system that LLMs need to interact with a SQL interpreter, reason, and organize the results. Yang et al. (2023) creates interactive code environments (Bash, SQL, and Python) based on static datasets (Yu et al., 2018; Lin et al., 2018; Austin et al., 2021) for LLM to act on. Huang et al. (2024) build a dataset of machine learning tasks that LLMs need to perform actions like reading/writing files, executing code. SWE-Bench (Jimenez et al., 2023) builds a dataset on real-world software engineering problems. Xu et al. (2024) introduces a benchmark that supports cross-environment GUI tasks over websites, desktop computers, or mobile phones.
@@
LLMs in Text-to-SQL Text-to-SQL benchmarks Yu et al. (2018); Li et al. (2023c) are proposed to test models on generating SQL queries given a question. Many works have been proposed to solve this task Askari et al. (2024); Talaei et al. (2024). Wang et al. (2024a) proposes a multi-agent framework, including a decomposer agent for Text-to-SQL generation and two auxiliary agents for tool execution and error refinement. Gao et al. (2023) provides a systematic review of prompt engineering for Text-to-SQL generation. C3-SQL (Dong et al., 2023), StructGPT Jiang et al. (2023), Din-SQL (Pourreza and Rafiei, 2024) propose frameworks targeting SQL generation that consist of several stages with different strategies such as self-consistency (Wang et al., 2022) or query decomposition. StateFlow (Wu et al., 2024b) introduces a framework with state and transitions to control the data exploration and selection in SQL tasks.
@@ 3.2 Question Generation
We want to create questions that can measure LLM investigation skill: to answer a question, the LLM agent needs to probe into the log data, analyze, and link related events to find the source of the malicious activity. While manual creation of these questions is expensive and does not scale, LLMs have proven useful in QA generation (Bhusal et al., 2024; Alam et al., 2024). A straightforward way is to ask an LLM to read all the details of an incident and draft Q&As. However, we found that this method produced generic questions that ignore the key concepts used to link the alerts, events, and entities together into a cohesive investigation. Moreover, these questions can lack a deterministic answer or ask about security knowledge not present in the database.
@@
4:  $V_{s}\leftarrow\textsc{Sample}(\textsc{GetFarthestEntities}(u_{s},u_{t}),k)$ $\triangleright$ Get start entities
5:  $v_{e}\leftarrow\textsc{Sample}(\textsc{GetFarthestEntities}(u_{t},u_{s}),1)$ $\triangleright$ Get answer
6:  $(q,a)\leftarrow M(P_{g},u_{s},V_{s},u_{t},v_{e})$ $\triangleright$ Generate question and answer
7:  $s\leftarrow M(P_{s},q,\textsc{ShortestPath}(u_{s},v_{e}))$ $\triangleright$ Generate solution path
@@ 3.3 Environment Setup
3:if $\text{check\_answer}(\mathcal{S}_{|\mathcal{S}|},\,a_{\text{sub}})$ then
4:  return 11 $\triangleright$ Perfect match on final answer
@@
6:  if $\text{check\_step}(\mathcal{S}_{i},\,a_{\text{sub}})$ then
7:   $r_{\text{sum}}\leftarrow r_{\text{sum}}+\mathit{discount}$ $\triangleright$ Add discounted reward for this step
8:  $\mathit{discount}\leftarrow\mathit{discount}\times\gamma$ $\triangleright$ Update discount for next older step
@@
To conduct a security investigation, the agent needs to interact with a given database. We set up a MYSQL docker environment following  Yang et al. (2023) to execute queries. In the environment, the agent can choose to output a query to be executed or submit the answer. We set a maximum number of steps to answer one question, so the solving process ends when the agent submits the answer or the maximum number of steps is reached. We set the maximum number of entries and the maximum character length that can be returned. Only when the returned content exceeds both the max character length and the max number of entries, do we cut the entries to satisfy the max number of entries.
By default, we employ LLM to evaluate the submitted answers (we use GPT-4o). Our evaluation consists of two steps (See Algorithm 2): 1. match the submitted answer $a_{sub}$ with the ground-truth. 2. If not correct, we check if the submitted answer contains any intermediate step solutions, which shows how much progress the agent makes and useful information it acquires. We assign a decayed reward starting from the last step. This reward not only checks the final answer but also performs a fine-grained evaluation of the agent’s intermediate steps. This also evaluates the agent’s ability to identify and extract key information (IoCs) from its investigation process. Since each step of the solution contains an unambiguous string, deterministic string comparison is allowed. Thus, the environment can provide reliable rewards for intermediate steps, which is a source of process supervision that is rarely available in other reasoning tasks such as math or coding (Lightman et al., 2023). This makes our environment particularly well-suited for future work on training agents with RL.
@@ 4.2 Comparison of Different Methods
Setup. To disentangle how different methods (e.g, prompting, test-time scaling) influence agent behavior, we evaluate six different configurations (Prompts in Figure 10 to 13). (1) Base The default prompt. (2) Strategy adds additional notes on using the alert tables as a reference to investigate, which is usually how a SOC analysts find information. (3) ReAct We follow ReAct Yao et al. (2022) to add 3 few-shot examples to base. These examples are selected from successful runs of a sampled train set. (4) Expel  Zhao et al. (2024) first distills rules from a small training split, then invokes this external memory, as well as retrieving similar examples during inference. (5) Best-of-N retries at most three times and returns the highest-reward trajectory. This is an oracle setting in which the agent needs the reward to determine if a problem needs to be rerun. (6) Reflection Shinn et al. (2023) extends Best-of-$N$ by letting the agent criticize its failed attempt, append the learned rule to the prompt, and retry. We run all methods on GPT-4o, GPT-4o-mini, and o3-mini (temp = 0, max_step = 15) and report average reward, interaction turns, and API cost in Table 3. (Also see a preliminary fine-tuning result on GPT-4o in Appendix E.2.)
@@ 5 Conclusion
Future Directions. Open-source models are catching up with propriety models in our environment, and they can be used for further distillation and training to boost performance. Our environment offers fine-grained process rewards that allow precise credit assignment across steps, which is an uncommon but valuable feature for our environment. Overall, this makes it a promising testbed for training LLM agents with reinforcement learning (Zhao et al., 2025). Another interesting direction is to utilize the inherent graph structure of the environment to define alternate state-action spaces for RL training.
@@ Appendix B Additional Background
  • Leveraging Cyber-security Domain Expertise: Applying deep knowledge of security frameworks and best practices (e.g., MITRE ATT&CK  [The MITRE Corporation, 2025b], CVE (Common Vulnerabilities and Exposures) [The MITRE Corporation, 2025a], OSINT (Open Source Intelligence) [Nordine, 2024], etc.) to guide investigation processes and validate findings. Drawing on historical context about common threat actor tactics and industry-specific threats to anticipate potential entry points or attack vectors.